I am really frustrated with this problem, can you help me?
5/x+7 + 4/x-3 = x-11/x^2+4x-21
My first step was to factor the equation x^2+4x-21 which I found to give me (x+7)(x-3). but now I don't know where to go from here. I would appreciate your assistance. Thank you

Answer by Boreal(15235)

Multiply the whole equation by x^2+4x-21, written as (x-3)(x+7) as you have.
The first term has the x+7 cancel
it is 5(x-3). The second term has the x-3 cancel and is 4(x+7). The answer is now x-11
Distribute, and you have 5x-15 + 4x +28=x-11
9x+13=x-11
8x=-24
x= -3
(5/4) -(4/6) = -14/24